Im working on solving linear equations by elimination.
whats 2x-y=9 and 3x+4y=-14

Answer by ReadingBoosters(3246)

Multiply the first equation by 4, to have the same y coefficents in differing signs, then add the equaitons and solve for x
4(2x-y=9) => 8x - 4y = 36
3x + 4y = -14
8x + 3x - 4y + 4y = 36 + -14
11x = 22
x = 2
...
Multiply the first equation by 3 and the second equation by -2, add the equations and solve for y
3(2x-y=9) ==> 6x - 3y = 27
-2(3x+4y=-14) ==> -6x - 8y = 28
6x + -6x + -3y + -8y = 27 + 28
-11y = 55
y = -5
...
Solution (2, -5)
……………….
Have more questions? Get them all answered at once.
HomeworkHelpers@readingboosters.com
Understand your homework, prepare for upcoming tests, improve your scores.
www.MyHomeworkAnswers.com